OpenStudy (sw33thearts):

Tom throws a ball into the air. The ball travels on a parabolic path represented by the equation h = -8t 2 + 40t, where h represents the height of the ball above the ground and t represents the time in seconds. How many seconds does it take the ball to reach its highest point? What ordered pair represents the highest point that the ball reaches as it travels through the air?

OpenStudy (sw33thearts):

please really need help

OpenStudy (aveline):

The highest point is the vertex. Do you know how to find the vertex?

OpenStudy (sw33thearts):

i kinda forgot through life

OpenStudy (aveline):

It's -b/2a Do you know what b and a are?

OpenStudy (sw33thearts):

I remeber the eqation but don't know b or a

OpenStudy (aveline):

h=at^2+bt+c This means that b is 40 and a is -8. (-40)/(-16)=2.5 This means that the x-value is 2.5 Now, to find the height, we plug in 2.5 for t: h=-8(2.5)^2+40(2.5) h=(-8)(6.25)+(100) h=-50+100 h=50 This means that after 2.5 seconds, the maximum height is 50 ft.

OpenStudy (sw33thearts):

thanks

OpenStudy (sw33thearts):

@Aveline